OpenStudy (anonymous):

Constructing a box. From a rectangular piece of cardboard having dimensions 20 inches x 30 inches, an open box is to be made by cutting out identical squares of area x^2 from each corner and turning up the sides. (a) Show that the volume of the box is given by the function V(x)=x(20-2x)(30-2x). (b) Find all positive values of x such that V(x)>0.

OpenStudy (anonymous):

height=x length=20-2(what u cut out)=20-2x width=30-2(what u cut out)=30-2x so v(x)=l*w*h=x(20-2x)(30-2x) b) (0,10)